OpenStudy (anonymous):

Solve the following system using substitution x+3y=8 2x-y=2 show work pls

OpenStudy (saifoo.khan):

x+3y=8 -1 2x-y=2 -2 Solving equation 1. x+3y=8 Making x the subject, x = 8 - 3y Now use this value of x and insert in equation 2. 2x - y = 2 2(8 -2y) - y = 2 Solve.

OpenStudy (anonymous):

That doesn't look like substitution e.e

OpenStudy (saifoo.khan):

That's exactly what a substitution method has. o_O

OpenStudy (anonymous):

So confused... Thanks lol

OpenStudy (saifoo.khan):

Where are you stuck? Feel free to throw your problems.